Four interior angles of a pentagon measure 88°, 118°, 132°, and 100°. What is the measure of the fifth interior angle?
Answer by ikleyn(52848)

.
Four interior angles of a pentagon measure 88°, 118°, 132°, and 100°. What is the measure of the fifth interior angle?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

The sum of interior angles of n-gon is (n-2)*180°.

For a pentagon it is (5-2)*180° = 3*180° = 540°.

Therefore, the 5-th interior angle is

540° - (88° + 118° + 132° + 100°) = 102°.

Answer. The missed 5-th angle is 102°.